# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ 2.1 The human-AI pipeline
We next discuss information a feasible policy can rely on. We assume that $\{\sigma_{k},\ell_{k},v_{k}\}_{k\in\mathcal{K}}$ are known; but $h_{k}$ is unknown initially and must be learned via samples from human reviewers, i.e., the data set $\mathcal{D}(t)$. Since there are many types of contents, we assume that a type $k$ comes with a known $d$-dimensional feature vector $\boldsymbol{\phi}_{k}\in\mathbb{R}^{d}$. The average cost satisfies a linear model $h_{k}=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star}$ where $\boldsymbol{\theta}^{\star}\in\mathbb{R}^{d}$ is an unknown vector. We assume that $\|\boldsymbol{\theta}^{\star}\|_{2},\|\boldsymbol{\phi}_{k}\|_{2}$ are bounded by a known value $U$, $|h_{k}|$ is bounded by a known value $H\geq 1$. We also define $\eta=\min(1,\min_{k\in\mathcal{K}}|h_{k}|)$ which is the margin of the average cost capped at $1$ (and can be 00). The platform has no information of $\{\lambda_{k}(t),N(t)\}_{t\in[T]}$. A policy is feasible if its decisions for any period $t$ are only based on the observed sample path $\{j(\tau),\mathcal{A}(\tau),\mathcal{S}(\tau)\}_{\tau<t}\cup\{j(t)\}$, the data set $\mathcal{D}(t)$ and the initial information $\{\sigma_{k},\ell_{k},v_{k},\boldsymbol{\phi}_{k}\}_{k\in\mathcal{K}}$.
@@ 2.2 Objectives and Benchmark
Recall that the total harmfulness of a post is $c_{j}v_{k}\sum_{t=t(j)}^{\min(t(j)+\ell_{k}-1,T)}O_{j}(t)$ and the optimal clairvoyant that knows $c_{j}$ will set $O^{\star}_{j}(t)=\mathbbm{1}\left(c_{j}\leq 0\right)$ for any $j$. Letting $\mathcal{J}(t)$ be the set of posts that arrive in the first $t$ periods, the loss of a policy $\pi$ with respect to this clairvoyant is thus:
$\mathcal{L}^{\pi}(T)\coloneqq\sum_{j\in\mathcal{J}(t)}c_{j}v_{k(j)}\sum_{t=t(j)}^{\min(T,t(j)+\ell_{k(j)}-1)}(O_{j}(t)-\mathbbm{1}\left(c_{j}\leq 0\right)).$
Due to the variance $\sigma_{k}$ in posts of the same type $k$, any post that is not reviewed incurs a positive loss in expectation even if the average $h_{k}$ is known. In addition, humans cannot review all posts due to limited capacity. As a result, aiming for vanishing loss is unattainable and we need to define a benchmark that captures both the effect of variance and capacity constraints. Our benchmark also needs to accommodate the non-stationarity in arrival rates $\{\lambda(t)\}$ and review capacity $\{N(t)\}$.
We consider a deterministic (fluid) benchmark where for every period $t$, there is a mass of $\lambda_{k}(t)$ posts from type $k$. The platform admits a mass of $a_{k}(t)$ posts to review and leaves a mass of $\lambda_{k}(t)-a_{k}(t)$ posts classified based on $h_{k}$ and not reviewed at all. Admitted posts receive human reviews immediately and thus incur no costs. For a non-admitted post $j$ of type $k$, the expected per-period loss of leaving this post on the platform is $r_{k}^{O}\coloneqq v_{k}\mathbb{E}\left[(c_{j})^{+}\,\middle|\,k(j)=k\right]$ and the loss of rejecting this post is $r_{k}^{R}\coloneqq v_{k}\mathbb{E}\left[-(c_{j})^{-}\,\middle|\,k(j)=k\right]$, where we denote $x^{+}=\max(x,0)$ and $x^{-}=\min(x,0)$. With the assumption of Gaussian noise, these quantities have explicit expressions (see Section 3). The per-period loss of an non-admitted post in the benchmark is thus $r_{k}\coloneqq\min(r_{k}^{O},r_{k}^{R})$, by rejecting a post if $h_{k}>0$ and keeping it if $h_{k}\leq 0$. Across its lifetime $\ell_{k}$ , each non-admitted type-$k$ post incurs loss $r_{k}\ell_{k}$. The expected total loss is then $\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}r_{k}\ell_{k}(\lambda_{k}(t)-a_{k}(t))$.
@@
$\textsc{Reg}^{\pi}(w,T)\coloneqq\mathbb{E}\left[\mathcal{L}^{\pi}(T)\right]/T-\mathcal{L}^{\star}(w,T)/T.$
@@ 3 Balancing Idiosyncrasy and Delay with Known Average Cost
Our starting point is the simpler setting where the average harm $h_{k}$ is known for every type and the number of types $K$ is small. In this setting, classification can be directly optimized by removing a post if and only if it has positive average harm, i.e., $Y_{k}(t)=\mathbbm{1}\left(h_{k}\leq 0\right)$. The two additional decisions (admission and scheduling) are not as straightforward and give rise to an interesting trade-off. To understand this trade-off, we decompose the loss of a policy $\mathcal{L}(\pi,T)$ into two components: idiosyncrasy loss and delay loss. In particular, for a period $t$, consider a new post $j$ of type $k$.
  • If the post is not admitted for review then it incurs an idiosyncrasy loss of $r_{k}\ell_{k}$ due to variance: it either stays on the platform for $\ell_{k}$ periods (if $h_{k}\leq 0$) and incurs a loss $r_{k}^{O}\ell_{k}=r_{k}\ell_{k}$, or it is removed from the platform for $\ell_{k}$ periods (if $h_{k}>0$) and incurs a loss $r_{k}^{R}\ell_{k}=r_{k}\ell_{k}$.
@@ 3.2 Balanced admission control for idiosyncrasy and delay (BACID)
In period $t$, a new post $j$ arrives with type $k$ and the algorithm removes it if $h_{k}>0$ and leaves it on the platform otherwise (Line 1). We admit this post to review if its idiosyncrasy loss $r_{k}\ell_{k}$, scaled by a hyper-parameter $\beta$ (Line 1), is greater than the current number of type $k$ posts in the review queue, $Q_{k}(t)$, i.e., $A_{k}(t)=\Lambda_{k}(t)\mathbbm{1}\left(\beta r_{k}\ell_{k}\geq Q_{k}(t)\right)$ (Line 1). This rule requires an explicit calculation of $r_{k}$. We give exact formulae of $r_{k}^{O},r_{k}^{R}$ as a function when the average cost is $h$:
$r_{k}^{O}(h)=v_{k}(h\Phi(h/\sigma_{k})+\sigma_{k}\varphi(h/\sigma_{k}))\qquad\text{and}\qquad r_{k}^{R}(h)=v_{k}(-h\Phi(-h/\sigma_{k})+\sigma_{k}\varphi(-h/\sigma_{k})).$ (5)
These functional forms (proved in Appendix B.2) provide flexibility when $h_{k}$ is unknown. We define $r_{k}(h)=\min(r_{k}^{R}(h),r_{k}^{O}(h))$ and use $r_{k}^{O},r_{k}^{R},r_{k}$ as shorthands for $r_{k}^{O}(h_{k}),r_{k}^{R}(h_{k}),r_{k}(h_{k})$, which are all known in this section because $h_{k}$ is known.
@@
Data: $h_{k},\ell_{k},\mu_{k}$
1 $\beta\leftarrow 1/\sqrt{K\ell_{\max}}$ // Admission Parameters
@@ 5       $r_{k}\leftarrow\min(r_{k}^{O}(h_{k}),r_{k}^{R}(h_{k}))$
8if the review finishes (with probability $N(t)\cdot\mu_{k}$) then  $\mathcal{S}(t)=M(t)$ else $\mathcal{S}(t)=\perp$
@@
Our main result is that BACID with $\beta=1/\sqrt{K\ell_{\max}}$ achieves an average regret of $O(w\sqrt{K\ell_{\max}})$.
@@
The above result is agnostic to the window size $w$; if $w$ is known, $\beta=\sqrt{w/(K\ell_{\max})}+K$ leads to a bound of $O(\sqrt{wK\ell_{\max}})$. Note that our benchmark becomes stronger as $w$ increases; the benchmark can schedule a post admitted in period $t$ for review in period $t+w$ without delay loss.
@@
Moreover, the policy admits a new post when $\beta r_{k}\ell_{k}\geq Q_{k}(t)$, which upper bounds the queue length by $\beta r_{k}\ell_{k}+1$. This implies a delay loss of $r_{\max}K(\beta r_{\max}\ell_{\max}+1)T$ where $r_{\max}$ is an upper bound over $r_{k}^{O}$ and $r_{k}^{R}$. Hence, a larger $\beta$ leads to more delay loss, which matches our intuition on the trade-off between idiosyncrasy and delay loss. Setting $\beta=1/\sqrt{K\ell_{\max}}$ balances this trade-off.
@@ 4 BACID with Learning: Optimism and Selective Sampling
We first restrict our attention to BACID admission rule: defer a post of type $k$ to human review if and only if $\beta r_{k}\ell_{k}\geq Q_{k}(t)$ where $r_{k}=\min(r_{k}^{O}(h_{k}),r_{k}^{R}(h_{k}))$ and, recalling Eq.(5):
$r_{k}^{O}(h)=v_{k}(h\Phi(h/\sigma_{k})+\sigma_{k}\varphi(h/\sigma_{k}))\qquad\text{and}\qquad r_{k}^{R}(h)=v_{k}(-h\Phi(-h/\sigma_{k})+\sigma_{k}\varphi(-h/\sigma_{k})).$
When the expected costs $h_{k}$ are unknown, we cannot directly compute $r_{k}$ and we need to instead use some estimate for $r_{k}$. A canonical way to resolve this problem in, e.g., bandits with knapsacks [AD19] is to use an optimistic estimate $\bar{r}_{k}(t)$. In particular, for each type $k$, we can compute the sample-average cost $\hat{h}_{k}(t)=\frac{\sum_{(j,c_{j})\in\mathcal{D}(t)\colon k(j)=k}c_{j}}{n_{k}(t)}$ where $n_{k}(t)=\sum_{(j,c_{j})\in\mathcal{D}(t)}\mathbbm{1}\left(k(j)=k\right)$ is the number of samples from type $k$; if $n_{k}(t)=0$, we can set $\hat{h}_{k}(t)=0$. We can then compute a confidence interval $[\underline{h}_{k}(t),\bar{h}_{k}(t)]$ which is, with high probability, valid, i.e., the true expected cost $h_{k}$ lies within the confidence interval (see Lemma 4.5):
$\underline{h}_{k}(t)=\max\left(-H,\hat{h}_{k}(t)-\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right),\bar{h}_{k}(t)=\min\left(H,\hat{h}_{k}(t)+\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right)$ (11)
where we recall that $H\geq 1$ is a known upper bound on the absolute value of $h_{k}$. We can then admit a post if and only if $\beta\bar{r}_{k}(t)\ell_{k}\geq Q_{k}(t)$ where $\bar{r}_{k}(t)=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h)$ is the most optimistic estimate for the per-period idiosyncrasy loss, i.e. it is the highest possible value that the per-period idiosyncrasy loss can have assuming that the sample-average cost is within the confidence interval. [^4] The solution $h$ giving the highest $r_{k}(h)$ has a simple form: if $0\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]$, the solution is zero; if $\bar{h}_{k}(t)<0$, the solution is $\bar{h}_{k}(t)$; otherwise, $\underline{h}_{k}(t)>0$, the solution is $\underline{h}_{k}(t).$ A formal proof is in Lemma C.1.
@@ 4.1 Why Optimism-Only is Insufficient for Classification
    classify a type-$k$ post as harmful if and only if $\hat{h}_{k}(t)\geq 0$ (similar to Line 1 in Algorithm 1);
@@
Such optimism-only heuristics generally work well in a constrained setting, such as bandits with knapsacks. The intuition is that assuming a valid confidence interval, we always admit a post that would have been admitted by BACID with known expected costs $h_{k}$. If we admit a post that would not have been admitted by BACID, we obtain one more sample; this shrinks the confidence interval which, in turn, limits the number of mistakes and leads to efficient learning.
Interestingly, this intuition does not carry over to our setting due to the additional error in the classification decisions of posts that are not admitted, for which the sign of $\hat{h}_{k}(t)$ may differ from that of $h_{k}$. To illustrate this point, consider the following instance with $K=2$ types of posts: texts (type-$1$) and videos (type-$2$); see (Figure 2 for the exact instance parameters. Texts have a much larger lifetime than videos ($\ell_{1}=49\ell_{2}$). Proposition 3 (proved in Appendix C.2) shows that BACID.UCB with $\beta\in(48/\ell_{1},1)$ and initial knowledge of $h_{1}$ does not review video posts at all with high probability and thus there is no video post in the dataset for the entire horizon, although there are $\Omega(T)$ arrivals of them. As a result, when classifying type-$2$ posts, since there is no data, the algorithm estimates $\hat{h}_{2}(t)=0$, and will remove all of videos, incurring $\Omega(\ell_{2})$ average regret. [^5] We assume removal when $\hat{h}_{k}(t)=0$. If the algorithm keep the post, the same issue exists by setting $h_{2}=1$.

### v3

@@ 2.1 The human-AI pipeline
We next discuss information a feasible policy can rely on. We assume that, for each type $k$, $\ell_{k}$ is known, but $\mathcal{F}_{k}$ is unknown initially and must be learned via samples from human reviewers, i.e., the data set $\mathcal{D}(t)$. The platform has no information of $\{\lambda_{k}(t),N(t)\}_{t\in[T]}$. We assume $\mathbb{E}_{k}\left[|c_{j}|\right]$ is bounded by a known value $r_{\max}\geq 1$ for any $k\in\mathcal{K}$ and that the cost distribution $\mathcal{F}_{k}$ is sub-Gaussian with a known variance proxy $\sigma_{\max}^{2}$ such that $\mathbb{E}_{k}\left[\exp(s(c_{j}-h_{k}))\right]\leq\exp(\sigma_{\max}^{2}s^{2}/2)$ for any $s\in\mathbb{R}$ and $k\in\mathcal{K}.$ A policy is feasible if its decisions for any period $t$ are only based on the observed sample path $\{j(\tau),\mathcal{A}(\tau),\mathcal{S}(\tau)\}_{\tau<t}\cup\{j(t)\}$, the data set $\mathcal{D}(t)$ and the initial information $r_{\max},\sigma_{\max}$ and $\{\ell_{k}\}_{k\in\mathcal{K}}$.
Remark 1.
Our model assumes a dual role for reviewers who conduct two tasks simultaneously: enforcement and labeling. On the one hand, reviewers decide whether to keep/remove a post; on the other hand, they also label the post for the AI. This is consistent with current practice  [^5] For example, in Meta’s content moderation pipeline, “Every time reviewers make a decision, we use that information to train our technology.”[Met22]. It is unclear whether conducting the two tasks separately may require different service times and lead to a better scheduling policy, which is an interesting future direction to explore.
@@ 2.2 Objectives and Benchmark
Recall that the total harmfulness of a post is $c_{j}\sum_{t=t(j)}^{\min(t(j)+\ell_{k}-1,T)}O_{j}(t)$ and the optimal clairvoyant that knows $c_{j}$ will set $O^{\star}_{j}(t)=\mathbbm{1}\left(c_{j}\leq 0\right)$ for any $j$. Letting $\mathcal{J}(t)$ be the set of posts that arrive in the first $t$ periods, the loss of a policy $\pi$ with respect to this clairvoyant is thus:
$\mathcal{L}^{\pi}(T)\coloneqq\sum_{j\in\mathcal{J}(t)}c_{j}\sum_{t=t(j)}^{\min(T,t(j)+\ell_{k(j)}-1)}(O_{j}(t)-\mathbbm{1}\left(c_{j}\leq 0\right)).$
Due to the uncertainty in the cost of a post, any post that is not reviewed incurs a positive loss in expectation even if the cost distribution is known. In addition, humans cannot review all posts due to limited capacity. As a result, aiming for vanishing loss is unattainable and we need to define a benchmark that captures both the effect of cost uncertainty and capacity constraints. Our benchmark also needs to accommodate the non-stationarity in arrival rates $\{\lambda(t)\}$ and review capacity $\{N(t)\}$.
We consider a deterministic (fluid) benchmark where for every period $t$, there is a mass of $\lambda_{k}(t)$ posts from type $k$. The platform admits a mass of $a_{k}(t)$ posts to review and leaves a mass of $\lambda_{k}(t)-a_{k}(t)$ posts classified based on $h_{k}$ and not reviewed at all. Admitted posts receive human reviews immediately and thus incur no costs. For a non-admitted post $j$ of type $k$, the expected per-period loss of keeping this post on the platform is $r_{k}^{O}\coloneqq\mathbb{E}_{k}\left[(c_{j})^{+}\right]$ and the loss of rejecting this post is $r_{k}^{R}\coloneqq\mathbb{E}_{k}\left[-(c_{j})^{-}\right]$, where we denote $x^{+}=\max(x,0)$ and $x^{-}=\min(x,0)$. The per-period loss of an non-admitted post in the benchmark is thus $r_{k}\coloneqq\min(r_{k}^{O},r_{k}^{R})$, by rejecting a post if $h_{k}>0$ or keeping it if $h_{k}\leq 0$ (this is because $r_{k}^{O}\leq r_{k}^{R}$ if and only if $r_{k}^{O}-r_{k}^{R}=h_{k}\leq 0$). Across its lifetime $\ell_{k}$ , each non-admitted type-$k$ post incurs loss $r_{k}\ell_{k}$. The expected total loss is then $\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}r_{k}\ell_{k}(\lambda_{k}(t)-a_{k}(t))$.
@@
$\textsc{Reg}^{\pi}(w,T)\coloneqq\mathbb{E}\left[\mathcal{L}^{\pi}(T)\right]/T-\mathcal{L}^{\star}(w,T)/T.$ (3)
@@ 3 Balancing Idiosyncrasy and Delay with Known Average Cost
Our starting point is the simpler setting where the cost distribution $\mathcal{F}_{k}$ is known for every type and the number of types $K$ is small. In this setting, classification can be directly optimized by keeping a post if and only if it has non-negative mean cost, i.e., $Y_{k}(t)=\mathbbm{1}\left(h_{k}\leq 0\right)$. The two additional decisions (admission and scheduling) are not as straightforward and give rise to an interesting trade-off. To understand this trade-off, we decompose the loss of a policy $\mathcal{L}(\pi,T)$ into two components: idiosyncrasy loss and delay loss. In particular, for a period $t$, consider a new post $j$ of type $k$.
  • If the post is not admitted for review, it incurs an idiosyncrasy loss of $r_{k}\ell_{k}$ due to cost uncertainty: it either stays on the platform for $\ell_{k}$ periods ($h_{k}\leq 0$) and incurs a loss $r_{k}^{O}\ell_{k}=r_{k}\ell_{k}$, or it is removed from the platform for $\ell_{k}$ periods ($h_{k}>0$) and incurs a loss $r_{k}^{R}\ell_{k}=r_{k}\ell_{k}$.
@@ 3.2 Balanced admission control for idiosyncrasy and delay (BACID)
In period $t$, a new post $j$ arrives with type $k$ and the algorithm removes it if $h_{k}>0$ and leaves it on the platform otherwise (Line 1). We admit this post to review if its idiosyncrasy loss $r_{k}\ell_{k}$, scaled by an admission parameter $\beta$, is greater than the current number of type $k$ posts in the review queue, $Q_{k}(t)$, i.e., $A_{k}(t)=\Lambda_{k}(t)\mathbbm{1}\left(\beta r_{k}\ell_{k}\geq Q_{k}(t)\right)$ (Line 1). This rule requires the value of $r_{k}$, which is known in this section because the cost distribution $\mathcal{F}_{k}$ is assumed known.
@@
Data: $h_{k},\ell_{k},\mu_{k}$     Admission parameter: $\beta\leftarrow 1/\sqrt{K\ell_{\max}}$
@@ 4       $r_{k}\leftarrow\min(r_{k}^{O},r_{k}^{R})$
7if the review finishes (with probability $N(t)\cdot\mu_{k}$) then  $\mathcal{S}(t)=M(t)$ else $\mathcal{S}(t)=\perp$
@@
Our main result is that BACID with $\beta=1/\sqrt{K\ell_{\max}}$ achieves an average regret of $O(w\sqrt{K\ell_{\max}})$, which improves the $\Omega(\ell_{\max})$ regret of static policies. In addition to this theoretical result, numerical results in Section 6.1 further show the efficiency of BACID.
@@
The above result is agnostic to the window size $w$; if $w$ is known, $\beta=\sqrt{w/(K\ell_{\max})}+K$ leads to a bound of $O(\sqrt{wK\ell_{\max}})$. Note that ($w$-fluid) becomes stronger as $w$ increases; this benchmark can schedule a post admitted in period $t$ for review in period $t+w$ without delay loss.
@@
Moreover, the policy admits a new post when $\beta r_{k}\ell_{k}\geq Q_{k}(t)$, which upper bounds the queue length by $\beta r_{k}\ell_{k}+1$. This implies a delay loss of $r_{\max}K(\beta r_{\max}\ell_{\max}+1)T$ where $r_{\max}$ is the known upper bound on $\mathbb{E}_{k}\left[|c_{j}|\right]$ and upper bounds $r_{k}^{O}$ and $r_{k}^{R}$. Hence, a larger $\beta$ leads to more delay loss, which matches our intuition on the trade-off between idiosyncrasy and delay loss. Setting $\beta=1/\sqrt{K\ell_{\max}}$ balances this trade-off.
@@ 4 BACID with Learning: Optimism and Selective Sampling
We first restrict our attention to BACID admission rule: defer a post of type $k$ to human review if and only if $\beta r_{k}\ell_{k}\geq Q_{k}(t)$ where $r_{k}=\min(r_{k}^{O},r_{k}^{R})$. When the cost distributions $\{\mathcal{F}_{k}\}_{k\in\mathcal{K}}$ are unknown, we cannot directly compute $r_{k}$ and we need to instead use some estimate for $r_{k}$. A canonical way to resolve this problem in, e.g., bandits with knapsacks [AD19] is to use an optimistic estimate $\bar{r}_{k}(t)$. In particular, for each type $k$, we can compute the sample-average estimation of $r_{k}^{O}$ and $r_{k}^{R}$ by $\hat{r}_{k}^{O}(t)=\frac{\sum_{(j,c_{j})\in\mathcal{D}(t)\colon k(j)=k}c^{+}_{j}}{n_{k}(t)}$ and $\hat{r}_{k}^{R}(t)=\frac{\sum_{(j,c_{j})\in\mathcal{D}(t)\colon k(j)=k}(-c^{-}_{j})}{n_{k}(t)}$ where $n_{k}(t)=\sum_{(j,c_{j})\in\mathcal{D}(t)}\mathbbm{1}\left(k(j)=k\right)$ is the number of samples from type $k$; if $n_{k}(t)=0$, we set $\hat{r}_{k}^{O}(t)=\hat{r}_{k}^{R}(t)=0$. The sample-average of the mean cost $h_{k}$ is given by $\hat{h}_{k}(t)=\hat{r}_{k}^{O}(t)-\hat{r}_{k}^{R}(t)$. We can then compute a confidence bound of $h_{k}$ and an upper confidence bound of $r_{k}$ by
$\underline{h}_{k}(t)=\max\left(-r_{\max},\hat{h}_{k}(t)-\sigma_{\max}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right),~{}\bar{h}_{k}(t)=\min\left(r_{\max},\hat{h}_{k}(t)+\sigma_{\max}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right)$ (11)
$\bar{r}_{k}(t)=\min\left(r_{\max},\min\left(\hat{r}_{k}^{O}(t),\hat{r}_{k}^{R}(t)\right)+4\sigma_{\max}\sqrt{\frac{\ln t}{n_{k}(t)}}\right)$ (12)
where we recall that $\sigma_{\max}^{2}$ is the known variance proxy of cost $c_{j}$ and $r_{\max}\geq 1$ is a known upper bound on $\mathbb{E}_{k}\left[|c_{j}|\right]$, thus an upper bound on $|h_{k}|,r_{k}^{O},r_{k}^{R},$ and $r_{k}$. We can then admit a post if and only if $\beta\bar{r}_{k}(t)\ell_{k}\geq Q_{k}(t)$.
@@ 4.1 Why Optimism-Only is Insufficient for Classification
    classify a type-$k$ post as harmful if and only if $\hat{h}_{k}(t)>0$ (similar to Line 1 in Algorithm 1);
@@
Such optimism-only heuristics generally work well in a constrained setting, such as bandits with knapsacks. The intuition is that assuming a valid upper confidence bound, we always admit a post that would have been admitted by BACID with known cost distributions $\{\mathcal{F}_{k}\}_{k\in\mathcal{K}}$. If we admit a post that would not have been admitted by BACID, we obtain one more sample; this shrinks the confidence interval which, in turn, limits the number of mistakes and leads to efficient learning.
Interestingly, this intuition does not carry over to our setting due to the additional error in the classification decisions of posts that are not admitted, for which the sign of $\hat{h}_{k}(t)$ may differ from that of $h_{k}$. To illustrate this point, consider the following instance with $K=2$ types of posts: texts (type-$1$) and videos (type-$2$). A text post has equal probability to be harmful or non-harmful (cost equal to ±$\pm 1$ with probability $0.5$) while a video post is harmful (cost equal to $1$) with probability $0.95$ and non-harmful (cost equal to $-1$) with probability $0.05$. Videos have a shorter lifetime than posts ($\ell_{2}=\ell_{1}/96$) and appear in the platform only after sufficient text posts are in the review system; see Figure 2 for the exact instance parameters.
Proposition 3 shows that BACID.UCB with $\beta\in(96/\ell_{1},1)$ and initial knowledge of $\mathcal{F}_{1}$ does not review video posts (for which $\mathcal{F}_{2}$ is unknown) at all with high probability and thus there is no video post in the dataset for the entire horizon, although there are $\Omega(T)$ arrivals of them. The proof, given in Appendix C.1, relies on the fact that the queue of type-$1$ posts is always longer than that of type-$2$ posts and thus no type-$2$ post will be scheduled for human review. As a result, when classifying type-$2$ posts, since there is no data, the algorithm estimates $\hat{h}_{2}(t)=0$, and keeps all videos even if they are likely harmful, incurring $\Omega(\ell_{2})$ average regret. [^6] We assume keeping the post when $\hat{h}_{k}(t)=0$. If the algorithm removes the post, the same issue exists by setting the cost distribution such that a type-$2$ post has $-1$ cost with probability $0.95$ and $+1$ cost with probability $0.05$. Our numerical simulation in Section 6.2 corroborates this theoretical observation.